OpenStudy (anonymous):

Algebra 1 help anyone?

OpenStudy (anonymous):

i took algebra 1 last year

OpenStudy (anonymous):

18x=18-18y 12x= 36- 18 Solve the system of linear equations seperate the x and y values

OpenStudy (anonymous):

please include all steps

OpenStudy (anonymous):

@erika_riera any idea?

OpenStudy (anonymous):

Uggggggggh It won't be that problem now my session just timed out

OpenStudy (anonymous):

opps srry was doing my history hw

OpenStudy (anonymous):

@tHe_FiZiCx99 pm me

OpenStudy (the_fizicx99):

Btw, I think you're missing a y in the second equation >.<

OpenStudy (anonymous):

try this website http://www.mathsisfun.com/algebra/linear-equations.html

OpenStudy (the_fizicx99):

Linear equations are easy, Convert them into y = mx + b and plot them, 18x=18-18y 18x = 18 - 18y +18y +18y 18y + 18x = 18 -18x -18x 18y = -18x + 18 /18 /18 /18 y = -x + 1 Now, in y = mx + b, M = the slope B = y intercept So, y = -x + 1 Plot the y intercept at (0,1) and go down 1, right one. It should look like this

OpenStudy (the_fizicx99):

Can you solve: 12x= 36- 18 ??